13.1.25 Adults and high school students were asked four trvia questions. The number of correct answers given by each participant was recorded To the right is a table showing the results Complete parts a) through e) ults otal a) If we select a participant at random, whats the probability we choose an adult who had one correct answer? 0089 (Round to three decimal places as needed) b) Among the participants who had one correct answer, what is the probability the person is an adult? 0650 (Round to three decimal places as needed) c) What is the probability that an adult participant had one correct answer? 017A (Round to three decimal places as needed) d) What is the probability that a student participant had no correct answers? 0 294 (Round to three decimal places as needed) e) What is the probability that a participant with no correct answers is a student? Question is complete. Tap on the red indicators to see incorrect answers All parts showing O Type here to search o@e??? ? 2 8 2 3 4 5 6 QW E R T Y U

media%2F749%2F749de305-c816-4e87-821b-be
Answer #1

a)P(adult with one correct answer) = count adult with one correct answer/total count

=34/315 = 0.108

b)P(adult/ one correct answer)

= P( adult and one correct answer) / P(one correct answer)

= (34/315) / ((34+15)/315)

= 0.694

c) P(one correct answer/ adult) = P( adult and one correct answer) / P(adult)

= (34/315) / (162/315)

= 0.210

d) P(student and no correct answer) = 41/315

= 0.130

e)P(student/ no correct answer) = P(student and no correct answer)/ P(no correct answer)

= (41/315) / (41+16)/315

= 0.719